OpenStudy (anonymous):

Find the equation for the line tangent to the graph of the given function at the indicated point. f(x)=x-x^2 at (-4,-20)

OpenStudy (anonymous):

f'(x) = 1- 2x -4 = 1- 2(-20) + b -5 = -40 + b b = 35 equation: y = 36 -2x might need to double check with someone else

OpenStudy (anonymous):

thanks for replying! the answer key says its y=9x+16. not sure how to get to that answer

OpenStudy (anonymous):

sorry i messed up unfortunately i forgotten my calc, all i know is that derviative will give you an equation tagent to the function but im not exactly sure how to do it for a specific point sorry

OpenStudy (anonymous):

no worries but thanks for trying

OpenStudy (campbell_st):

ok...do you understand what the 1st derivative gives you..?

OpenStudy (anonymous):

1-2x?

OpenStudy (campbell_st):

the 1st derivative gives the general equation of the slope of the tangent. f'(x) = 1 - 2x you need to find the slope at the point where x = -4 so evaluate f'(-4) = 2 - 2(-4)

OpenStudy (anonymous):

=10

OpenStudy (campbell_st):

ok... so the slope of the tangent is m = 10 so you have a point (-4, -20) and a slope m = 10 can you find the equation of the line..?

OpenStudy (campbell_st):

but you need to check you solution for the value of the slope... f'(-4) = 1 - 2(-4)

OpenStudy (anonymous):

yay i got it! so the slope is 9 y-(-20)=9(x-(-4) y+20=9x+36 -20 -20 y=9x+16 thanks!!!!! :)

OpenStudy (campbell_st):

that correct... just remember the derivative is the equation of the slope of the tangent... substitute the x value to get the slope at a specific point... glad it helped.